OpenStudy (anonymous):

Differentiate both sides of the given equation directly. If y^(1/2)=x/[(xy^(1/2)+y], find dy/dx.

OpenStudy (anonymous):

Implicit differentiation? Were you able to start the problem?

OpenStudy (anonymous):

You want to treat y as a function of x... so d/dx of y is y' and d/dx of some function of y (like y^2 for example) requires the chain rule: d/dx (y^2) = 2*y*y'

OpenStudy (anonymous):

(((([xy^(1/2)+y]^2/2y^(1/2) + x{x/[2y^(1/2) +1]} ))))(dy/dx) = y i dont know what to do next.

OpenStudy (anonymous):

trying to read that...

OpenStudy (anonymous):



OpenStudy (anonymous):

ans is 2(1-y)/[2x+3y^(1/2)] i have no idea how to do this.

OpenStudy (anonymous):

well, it looks like you're on the right track... just divide by the coeffecient of dy/dx and you should be done.

OpenStudy (anonymous):

It's almost impossible to read what you've got there, so I'll try to rewrite it... tell me if I'm right.

OpenStudy (anonymous):